OpenStudy (123ab456c):

What is the area in square centimeters of a rectangle with a width of 60.4 millimeters, and a length of 85.7 millimeters? A. 51.7628 cm2 B. 517.628 cm2 C. 5176.28 cm2 D. 51,762.8 cm2

OpenStudy (alexandervonhumboldt2):

area=l*w

OpenStudy (alexandervonhumboldt2):

area=60.4*85.7=?

OpenStudy (alexandervonhumboldt2):

and 1 cm=10 mills

OpenStudy (alexandervonhumboldt2):

so you'll need to convrert

OpenStudy (123ab456c):

is the answer C

OpenStudy (alexandervonhumboldt2):

no

OpenStudy (alexandervonhumboldt2):

you forgot to convert form millimeters to centimeters

OpenStudy (123ab456c):

IS IT D then

OpenStudy (alexandervonhumboldt2):

no

OpenStudy (alexandervonhumboldt2):

pleae show your work

OpenStudy (123ab456c):

why

OpenStudy (alexandervonhumboldt2):

openstudy is for learning, not for answers

OpenStudy (123ab456c):

thats why i am here to get the ANSWER if i wanted you to explain it i woulve asked you!!!:(

OpenStudy (alexandervonhumboldt2):

ok here are all answers to your test http://pages.citebite.com/a4o3b6r4a2tjy

OpenStudy (123ab456c):

>:)

OpenStudy (123ab456c):

i really dont care becuase erybody I ASK TELLS ME THE ANSWER !!!!

OpenStudy (alexandervonhumboldt2):

@Jamierox4ev3r

OpenStudy (alexandervonhumboldt2):

all those who give answers are violating rules.

OpenStudy (123ab456c):

WHO CARES EVERYONE DOES IT!!! ECCEPT FOY YOU AND POOJA!!!!

OpenStudy (alexandervonhumboldt2):

no

OpenStudy (alexandervonhumboldt2):

you gonna see that 80% of openstudy

OpenStudy (123ab456c):

LIES!!! YOU ARENT THE ONE ASKING QUESTIONS SO YOU DONT KNOW!!!! DONT QUESTION ME!

OpenStudy (123ab456c):

WILL someone please tell me the answer

OpenStudy (123ab456c):

why cant you tell me the answer and then EXPLAIN it can you PLEASEEEEE do that

OpenStudy (alexandervonhumboldt2):

i explained, it is not hard o mutlply 85.7*60.4 and then divide by 10

OpenStudy (nathalie0918):

@AlexandervonHumboldt2 is right people cant just jive you the answer because this site is about learning, you have to come up with an answer your self with helpers to guide you

OpenStudy (123ab456c):

you to forget it!!

OpenStudy (123ab456c):

the answer is FREKIN B

OpenStudy (alexandervonhumboldt2):

here is no answer if you don't know how to do it. :)